OpenStudy (anonymous):

Approximately how many degrees are in the measure of an interior angle of a regular seven sided polygon? 102.9 128.6 180.0 231.4

OpenStudy (anas.p):

To find the sum of interior angles, the formula is \[(2n-4)\times90\] here n=7 so \[((2\times 7)-4)\times90\]\[10\times 90\]900 is the sum of interior angles. \[900\div7\] because no. of sides. 128.57 Answer B